sara has just renter her first apartment strating december 1 before beginning college in january. The apartment has a washer and a dryer hookup, so sara wanted to buy the appliances to avoid trips to the laundomat. The saturday newspaper had an advertisement for a local appliance store offering 90 days same as cash financing sara asked how the financing worked and learned that she could pay for the washer and dryer anytime during the first 90 days for the sales purchase price plus sales tax. If she waited longer, she would have to pay the purchase price, plus sales tax, plus 26.8% annual simple interest for the first 90 days plus 2% simple interest per month on unpaid balance after 90 days with minimum payments of $50 per month after 90 days period. Together the washer and dryer cost 699 plus the 8.25% sales tax. sara knew that her tax refund from the irs would be 1000 so she bought the washer and dryer confident that she could pay off the balance within the 90 days.
1. if sara psys off the balance within 90 days how much will she pay?
2. If sara bought the washer and dryer on December 15 using the exact interest, what is her deadline for paying no interest in a non leap year? in a leap year will the finance company be more likely to use or ordinary interest and why?
3. If sara irs refund does not come until april 1 what is her payoff amount assume ordinary interest and a non leap year
4. How much did it cost her to pay off this loan 17 days late? what annual simple interet rate does this amonut to?

Answer by drk(1908)

I think I am on the right track here . . .
Part 1:pay off within 90 days means purchase price plus sales tax. 699 * (1.0825) = $756.67
Part 2: Buys december 15. 90 days in non leap year = March 15; 90 days in leap year = March 16. The finance company I think will use ordinary interest.
part 3: If she pays on April 1st, that is 16 days late. She pays 699 + 8.25% sales tax + .268(90/365) + .02*699*1 = $770.71
part 4: Compared to payoff within 90 days, it cost her $14.04. This is 14.04/699 = 2%.